If Wild Widgets, Inc., were an all-equity company, it would have a beta of .90. The company has a target debt-equity ratio of .75. The expected return on the market portfolio is 11 percent and Treasury bills currently yield 3.6 percent. The company has one bond issue outstanding that matures in 20 years, a par value of $1,000, and a coupon rate of 6.5 percent. The bond currently sells for $1,075. The corporate tax rate is 23 percent.

  

a.

What is the company’s cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the company’s cost of equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What is the company’s weighted average cost of capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

CASE 1: ASSUMING ANNUAL COUPON PAYMENTS
1.
=RATE(20,6.5%*1000,-1075,1000)*(1-23%)
=4.50747%

2.
=3.6%+0.90*(1+(1-23%)*0.75)*(11%-3.6%)
=14.10615%

3.
=0.75/1.75*(RATE(20,6.5%*1000,-1075,1000)*(1-23%))+1/1.75*(3.6%+0.90*(1+(1-23%)*0.75)*(11%-3.6%))
=9.99243%

CASE 2: ASSUMING SEMI-ANNUAL COUPON PAYMENTS
1.
=RATE(20*2,6.5%*1000/2,-1075,1000)*2*(1-23%)
=4.51102%

2.
=3.6%+0.90*(1+(1-23%)*0.75)*(11%-3.6%)
=14.10615%

3.
=0.75/1.75*(RATE(20*2,6.5%*1000/2,-1075,1000)*2*(1-23%))+1/1.75*(3.6%+0.90*(1+(1-23%)*0.75)*(11%-3.6%))
=9.99395%